# Massive deformations of supersymmetric Yang-Mills matrix models

### v1

@@ 7 Three-dimensional matrix model
In appendix A.4, we find a unique mass-deformation of the $D=3$ pure SYM matrix model that carries two supersymmetries together with $SU(N)$ gauge invariance. This model is parameterized by two mass-parameters $\mu_{1},\mu_{2}$. The action reads
$\begin{split}S_{D=3}=\frac{1}{g_{\mathrm{YM}}^{2}}\operatorname{Tr}&\Biggl{[}-\frac{1}{4}[X_{i},X_{j}]^{2}-\frac{i}{2}\bar{\psi}\Gamma^{i}[X_{i},\psi]\\ &+\mu_{1}\mu_{2}X_{i}X_{i}+i\frac{2}{3}(\mu_{1}+\mu_{2})\epsilon_{ijk}X_{i}X_{j}X_{k}+i\mu_{1}\bar{\psi}\psi\Biggr{]}\,,\end{split}$ (74)
@@
In Lorentzian signature, the bosonic symmetry group would be $SO(2,1)$, and there are 2 supercharges. In appendix A.4, we obtain that the superalgebra is gauged $OSp(1|2)$.
@@
Finally, let us mention observables in dimensionless units $(g_{\mathrm{YM}}\to 1,\mu_{i}\to\Omega_{i}\equiv\mu_{i}/\sqrt{g_{\mathrm{YM}}})$. In the strict massless limit at fixed $N$ and fixed ratio $\Omega_{1}/\Omega_{2}$, we can integrate out off-diagonals as shown in appendix C. This allows to compute a large class of observables, including e.g.
@@ 8 Discussion
In this work we have presented a complete classification of mass deformations of the SYM matrix models in various $D$ that keep the number of supersymmetries unchanged and retain gauge invariance, as summarized in table 1. This analysis establishes the uniqueness of the polarized IKKT model. In $D=4$ we identified two deformations with a positive semi-definite Pfaffian. The first, type I, serves as a useful benchmark, featuring a single mass parameter and only a trivial saddle, while the second, type II, has a richer structure with two mass parameters and a dominating fuzzy-sphere saddle. Numerical simulations of these toy models would be of considerable interest.
@@
Moreover, we derived only a partial set of analytic expressions for these new models. A full supersymmetric localization treatment along the lines of Pestun:2007rz ; Asano:2012zt ; Komatsu:2024ydh remains to be developed. In particular, one could compute the partition functions, as well as protected observables of the form $\langle f(X+iY)\rangle$ where $X$ and $Y$ are two distinct bosonic matrices among $X_{I=1,\ldots,D}$. For example, assume $f$ is a homogeneous polynomial in $X+iY$ of degree $k\neq 0$. This would be an interesting observable when the $SO(2)$ symmetry which rotates $X$ and $Y$ is broken by the mass-deformation. Otherwise, these expectation values would vanish identically as they carry a charge $k\neq 0$ under $SO(2)$. Therefore, in the $D=4$ type I model and the $D=3$ model where $SO(D)$ is preserved, only the partition function can be computed thanks to supersymmetric localization. However, in the other models which all break $SO(D)$, it would be interesting to compute e.g. the observables $\langle\mathrm{Tr}(X_{1}+iX_{D})^{2n}\rangle$ which would yield a non-trivial result.
@@
Finally, let us mention that our motivation for this work was based on finding a variety of toy models that would mimic features of polarized IKKT. However, the matrix models we constructed may have physical relevance in their own right, or at least describe some sector of models of physical interest. Both are true in the case of massive SYM quantum mechanics, as explained in Kim:2006wg . These quantum mechanical models were argued to be dual to non-critical M-theory on curved backgrounds. Moreover, some of them could be obtained from M5-brane embeddings in the 11-dimensional pp-wave geometry, or from compactifying four-dimensional SYM on a three-sphere. [^21] BMN arises from a sector of $\mathcal{N}=4$ SYM while some of the other massive SYM quantum mechanical models arise from four-dimensional $\mathcal{N}=2$ and $\mathcal{N}=1$. It would be interesting to explore whether there exist analogous physical setups which are dual or reduce to our matrix models.
@@ Acknowledgements
We thank S. Komatsu, J. Matos, H. Murali, J. Oliveira, J. Penedones, P. Vieira, A. Vuignier, T. Wiseman (via J. Penedones), X. Zhao for helpful discussions. This work was supported by the Simons Foundation grant 488649 (Simons Collaboration on the Nonperturbative Bootstrap) and by the Swiss National Science Foundation through the project 200020_197160 and through the National Centre of Competence in Research SwissMAP.
@@ A.1 Uniqueness proof of polarized IKKT
We’re thus left with $S_{IJ}$ and $m_{IJK}$ as unknowns. We now consider the second set of constraints (25). The structure is analogous to before. Thus, we get
@@
Plugging in $F^{I}$ in terms of $M$ and obtain that $\tilde{E}_{1}$, $\tilde{E}_{2}$, and $\tilde{E}_{3}$ read respectively
@@
$\begin{split}S=\frac{1}{g_{\mathrm{YM}}^{2}}\operatorname{Tr}&\Biggl{[}-\frac{1}{4}[X_{I},X_{J}]^{2}-\frac{i}{2}\bar{\psi}\Gamma^{I}[X_{I},\psi]\\ &+\frac{3\mu^{2}}{4^{3}}X_{i}X_{i}+\frac{\mu^{2}}{4^{3}}X_{p}X_{p}+i\frac{\mu}{3}\epsilon^{ijk}X_{i}X_{j}X_{k}-\frac{1}{8}\mu\bar{\psi}\Gamma^{123}\psi\Biggr{]}\,,\end{split}$ (104)
@@
In (Komatsu:2024ydh, , App.C.1), we found after introducing a set of auxiliary fields that the anticommutator of two supercharges yields a sum of $SO(3)$ and $SO(7)$ generators. Thus, in Lorentzian signature, the superalgebra is either $F(4)$ (containing $SO(2,1)\times SO(7)$) or $F^{1}(4)$ (containing $SO(3)\times SO(6,1)$) VanProeyen:1999ni .
@@ A.2 Six-dimensional classification
$\begin{split}S=S_{0}+\frac{1}{g_{\mathrm{YM}}^{2}}\mathrm{Tr}\Biggl{\{}\mu\epsilon^{ab}\bar{\psi}_{b}M\psi_{a}+\mu\bar{\psi}_{a}M^{ab}\psi_{b}+\mu S_{IJK}X^{I}X^{J}X^{K}+\mu^{2}S_{IJ}X^{IJ}\Biggr{\}}\,,\end{split}$ (108)
@@
$\begin{split}\delta_{\epsilon}X^{I}&=\epsilon^{ab}\bar{\epsilon}_{b}\Gamma^{I}\psi_{a}\,,\\ \delta_{\epsilon}\psi_{a}&=\frac{i}{2}\Gamma^{IJ}\epsilon_{a}[X_{I},X_{J}]+\mu F^{I}\epsilon_{a}X_{I}+\mu F^{I}_{ab}\epsilon_{b}X_{I}\,,\end{split}$ (109)
@@
In order for the Myers term to be Hermitian, $S_{IJK}$ must be purely imaginary. [^26] Note that if we don’t impose Hermicity of the Myers term, there is a solution where everything vanishes except the Myers term and $f_{IJK}^{A}$. This constrains
@@
From these we obtain [^27] Note that because of the second line, the fully antisymmetrized part over $I,J,K$ of $f^{I}_{abJK}$ has to vanish. Thus contrarily to the previous case there’s no free antisymmetric anti-self-dual part.
@@
$\begin{split}\Bigl{(}&-\epsilon^{cb}m_{JKN}f_{ca}^{I}+6\epsilon^{cb}m_{L[JK}f_{caN]L}^{I}-m_{[J}^{ab}f_{KN]}^{I}-m_{[J}^{cb}f_{caKN]}^{I}\\ &+6\epsilon^{ab}m_{L[JK}f_{N]L}^{I}\Bigr{)}\Gamma^{JKN}(\mathbb{1}+\Gamma_{*})=0\,,\end{split}$ (126)
@@ A.2.1 Type I : $SO(6)\to SO(3)\times SO(3)$
[…]
@@
Now taking a linear combination of these two equations [^29] More specifically, $E^{abN}_{JKN}-\frac{i}{3!}E^{abN}_{PQR}\epsilon^{PQRJKN}-F_{JK}^{ab}$, one obtains that the only possibility is $m^{ab}_{I}=0$. For convenience we choose $m=-1/96$. Putting everything together, we found the first solution
$\begin{split}S=\frac{1}{g_{\mathrm{YM}}^{2}}\operatorname{Tr}&\Biggl{[}-\frac{1}{4}[X_{I},X_{J}]^{2}-\frac{i}{2}\epsilon^{ab}\bar{\psi}_{b}\Gamma^{I}[X_{I},\psi_{a}]-\frac{1}{8}\mu\epsilon^{ab}\bar{\psi}_{b}\Gamma^{123}\psi_{a}\\ &+i\frac{\mu}{3}\epsilon_{ijk}X_{i}X_{j}X_{k}+\frac{3\mu^{2}}{2^{6}}X_{i}X_{i}+\frac{\mu^{2}}{2^{6}}X_{p}X_{p}\Biggr{]}\,,\end{split}$ (133)
@@
Note that this action is invariant under $SO(2,1)\times SO(3)$ rotations, and under $SU(2)$ transformations acting on the symplectic index, $\psi_{a}\to U_{ab}\psi_{b},U\in SU(2)$. The total bosonic algebra is thus isomorphic to $SO(4)\times Sl(2,\mathbb{R})$ which is the bosonic algebra of the exceptional continuous family of superalgebras $D(2,1,\alpha)$. To ensure that $D(2,1,\alpha)$ is indeed the superalgebra and compute $\alpha$, it is important to compute the anticommutators of two supercharges. To obtain a superalgebra which closes off-shell, we will introduce a set of auxiliary fields $K_{m=1,\ldots,3}$, and consider the action
@@
$\begin{split}\delta_{\epsilon}X^{I}&=\bar{\epsilon}\Gamma^{I}\psi\,,\\ \delta_{\epsilon}\psi_{a}&=\frac{i}{2}\Gamma^{IJ}\epsilon_{a}[X_{I},X_{J}]+\frac{3\mu}{8}\Gamma^{123}\Gamma^{i}\epsilon_{a}X_{i}+\frac{\mu}{8}\Gamma^{123}\Gamma^{p}\epsilon_{a}X_{p}+i\nu_{a}^{m}K_{m}\,,\\ \delta_{\epsilon}K_{m}&=-\epsilon^{ab}\bar{\nu}_{b}^{m}\Gamma^{I}[X_{I},\psi_{a}]+i\frac{\mu}{4}\epsilon^{ab}\bar{\nu}_{b}^{m}\Gamma^{123}\psi_{a}\,,\end{split}$ (136)
where $\nu_{a}^{m}(\epsilon)$ is a spinorial parameter that will be constrained in terms of $\epsilon$ to obtain closure off-shell. This construction is inspired from Berkovits:1993hx ; Asano:2012zt . Now we can compute the action of the anticommutator $\{\delta_{\epsilon},\delta_{\eta}\}$ on the relevant matrices exactly as in (Komatsu:2024ydh, , App.A.1). Imposing the closure of the algebra with the introduction of the auxiliary fields imposes conditions on $\nu_{m}^{a}$ that restricts it in terms of $\epsilon$. To obtain these constraints, it is sufficient to look at the undeformed model ($\mu=0$). We obtain in particular
@@
$\begin{split}\{Q_{\alpha a},Q_{\beta b}\}&=2i\epsilon^{ab}(\mathcal{C}\Gamma^{I})_{\alpha\beta}[X_{I},\cdot]-i\frac{3\mu}{8}\epsilon_{ab}(\mathcal{C}\Gamma^{123}\Gamma^{ij})_{\alpha\beta}M_{ij}\\ &\ \ +i\frac{\mu}{8}\epsilon_{ab}(\mathcal{C}\Gamma^{123}\Gamma^{pq})_{\alpha\beta}M_{pq}+\frac{i\mu}{2}(\mathcal{C}\Gamma^{123})_{\alpha\beta}T_{ab}\,,\end{split}$ (140)
@@
$\begin{gathered}\{Q_{\alpha^{\prime}\alpha^{\prime\prime}a},Q_{\beta^{\prime}\beta^{\prime\prime}b}\}=-i\frac{3\mu}{4}\epsilon_{ab}\epsilon_{\alpha^{\prime}\beta^{\prime}}T^{(1)}_{\alpha^{\prime\prime}\beta^{\prime\prime}}+i\frac{\mu}{4}\epsilon_{ab}\epsilon_{\alpha^{\prime\prime}\beta^{\prime\prime}}T^{(2)}_{\alpha^{\prime}\beta^{\prime}}+i\frac{\mu}{2}\epsilon_{\alpha^{\prime}\beta^{\prime}}\epsilon_{\alpha^{\prime\prime}\beta^{\prime\prime}}T_{ab}\,,\\ [T_{ab},Q_{\alpha^{\prime}\alpha^{\prime\prime}c}]=-\frac{i}{2}(\epsilon_{ac}Q_{\alpha^{\prime}\alpha^{\prime\prime}b}+\epsilon_{bc}Q_{\alpha^{\prime}\alpha^{\prime\prime}a})\,,\\ [T^{(1)}_{\alpha^{\prime\prime}\beta^{\prime\prime}},Q_{\alpha^{\prime}\gamma^{\prime\prime}a}]=-\frac{i}{2}(\epsilon_{\alpha^{\prime\prime}\gamma^{\prime\prime}}Q_{\alpha^{\prime}\beta^{\prime\prime}a}+\epsilon_{\beta^{\prime\prime}\gamma^{\prime\prime}}Q_{\alpha^{\prime}\alpha^{\prime\prime}a})\,,\\ [T^{(2)}_{\alpha^{\prime}\beta^{\prime}},Q_{\gamma^{\prime}\alpha^{\prime\prime}a}]=-\frac{i}{2}(\epsilon_{\alpha^{\prime}\gamma^{\prime}}Q_{\beta^{\prime}\alpha^{\prime\prime}a}+\epsilon_{\beta^{\prime}\gamma^{\prime}}Q_{\alpha^{\prime}\alpha^{\prime\prime}a})\,,\\ [T_{ab},T_{cd}]=-i(\epsilon_{ab}T_{bd}+\epsilon_{bd}T_{ac})\,,\end{gathered}$ (149)
where the last line also applies to $T^{(1)}$ and $T^{(2)}$. As a consistency check, note that the three coefficients in the anticommutators sum to zero, $-\frac{3}{4}+\frac{1}{4}+\frac{1}{2}=0$ as required by Jacobi identities Frappat:1996pb . Up to rescaling the supercharge, this is exactly the $D(1,2,\alpha=\frac{1}{2})$ superalgebra that can be found in the literature Ivanov:2002pc ; Frappat:1996pb .

### v2

@@ 7 Three-dimensional matrix model
In appendix A.4, we find a unique mass deformation of the $D=3$ pure SYM matrix model that carries two supersymmetries together with $SU(N)$ gauge invariance. This model is parameterized by two mass parameters $\mu_{1},\mu_{2}$. The action reads
@@
where $i=1,2,3$. The deformation preserves the global $SO(3)$ symmetry. For the bosonic quadratic term to be real and positive, we restrict to $\mu_{i}\in\mathbb{R},\mu_{1}\mu_{2}>0$ [^26] We discard the case $\mu_{1}=-\mu_{2}=i\mu$ with $\mu\in\mathbb{R}$ because the Pfaffian would be not only sign-changing, but also complex.. The two supersymmetries get deformed to
@@
In Lorentzian signature, the bosonic symmetry group would be $SO(2,1)$, and there are 2 supercharges. In appendix A.4, we obtain that the superalgebra is gauged $OSp(1|2)$. The one-parameter subfamily $\mu_{2}=3\mu_{1}$ can be obtained from SYM on $S^{n}$ with 2 supersymmetries, extrapolating Blau:2000xg to n=0n=0.272727In the notation of Blau:2000xg, it arises from the family B, with $(d=3,n=0)$
@@
Finally, let us mention observables in dimensionless units $(g_{\mathrm{YM}}\to 1,\mu_{i}\to\Omega_{i}\equiv\mu_{i}/\sqrt{g_{\mathrm{YM}}})$. In the strict massless limit at fixed $N$ and fixed ratio $\Omega_{1}/\Omega_{2}$, we can integrate out off-diagonals as shown in appendix C. This allows us to compute a large class of observables, including e.g.
@@ 8 Discussion
In this work we have presented a complete classification of mass deformations of the SYM matrix models in dimensions $D=3,4,6,10$ that keep the number of supersymmetries unchanged and retain gauge invariance, as summarized in table 1. This analysis establishes the uniqueness of the polarized IKKT model. In $D=4$ we identified two deformations with a positive semi-definite Pfaffian. The first, type I, serves as a useful benchmark, featuring a single mass parameter and only a trivial saddle, while the second, type II, has a richer structure with two mass parameters and a dominant fuzzy-sphere saddle. Below we provide a list of interesting future directions.
  • The main motivation for this work was to construct numerically tractable toy models. It would be very interesting to explore whether bootstrap or Monte Carlo methods can be efficiently applied to extract new, non-perturbative information. In particular, one interesting target would be to determine whether the massive SYM matrix models become commutative at strong ’t Hooft coupling ($N\to\infty$, $\lambda\equiv N/\Omega^{4}\gg 1$). If this is the case, one should be able to show numerically that $\langle\mathrm{Tr}[X_{I},X_{J}]^{2}\rangle/\langle\mathrm{Tr}\{X_{I},X_{J}\}^{2}\rangle$ becomes very small in this regime.
@@
  • In this paper, we derived only a partial set of analytic expressions for the new models. A full supersymmetric localization treatment along the lines of Pestun:2007rz; Asano:2012zt; Komatsu:2024ydh remains to be developed. In particular, one could compute the partition functions, as well as protected observables of the form $\langle f(X+iY)\rangle$ where $X$ and $Y$ are two distinct bosonic matrices among $X_{I=1,\ldots,D}$. For example, assume $f$ is a homogeneous polynomial in $X+iY$ of degree $k\neq 0$. This quantity would be an interesting observable when the $SO(2)$ symmetry that rotates $X$ and $Y$ is broken by the mass deformation. Otherwise, these expectation values would vanish identically as they carry a charge $k\neq 0$ under $SO(2)$.
@@
  • The motivation for this work was to find a variety of toy models that mimic features of the polarized IKKT model. However, the matrix models we constructed may have relevance in their own right, or at least describe some sector of models of interest. Both are true in the case of massive SYM quantum mechanics, as explained in Kim:2006wg. These quantum-mechanical models were argued to be dual to non-critical M-theory on curved backgrounds. Moreover, some of them could be obtained from M5-brane embeddings in the 11-dimensional pp-wave geometry, or from compactifying four-dimensional SYM on a three-sphere. [^28] BMN arises from a sector of $\mathcal{N}=4$ SYM while some of the other massive SYM quantum mechanical models arise from four-dimensional $\mathcal{N}=2$ and $\mathcal{N}=1$. It would be interesting to explore whether there exist analogous setups that are dual to, or reduce to, our matrix models.
  • A subset of the matrix models could be obtained from the $n=0$ extrapolation of SYM theories on $S^{n}$ Blau:2000xg. However, only one-parameter families can be extracted in this way, so the full two-parameter families of the $D=3$ and $D=4$ type II models cannot be derived. Moreover, the $D=4$ type I model cannot be obtained in such a way. It would be interesting to explore whether one can define SYM theories on other higher-dimensional manifolds, whose extrapolation to zero dimension reduces to the remaining matrix models.
@@ Acknowledgements
We thank M. Blau, N. Bobev, S. Komatsu, J. Matos, H. Murali, J. Nishimura, J. Oliveira, J. Penedones, P. Vieira, A. Vuignier, T. Wiseman (via J. Penedones), X. Zhao for helpful discussions. This work was supported by the Simons Foundation grant 488649 (Simons Collaboration on the Nonperturbative Bootstrap) and by the Swiss National Science Foundation through the project 200020_197160 and through the National Centre of Competence in Research SwissMAP.
@@ A.1 Uniqueness proof of polarized IKKT
We are thus left with $S_{IJ}$ and $m_{IJK}$ as unknowns. We now consider the second set of constraints (25). The structure is analogous to before. Thus, we get
@@
Plugging in $F^{I}$ in terms of $M$, we obtain that $\tilde{E}_{1}$, $\tilde{E}_{2}$, and $\tilde{E}_{3}$ read respectively
@@
[…]
@@
In (Komatsu:2024ydh, App.C.1), we found after introducing a set of auxiliary fields that the anticommutator of two supercharges yields a sum of $SO(3)$ and $SO(7)$ generators. Thus, in Lorentzian signature, the superalgebra is either $F(4)$ (containing $SO(2,1)\times SO(7)$) or $F^{1}(4)$ (containing $SO(3)\times SO(6,1)$) VanProeyen:1999ni.
@@ A.2 Six-dimensional classification
[…]
@@
[…]
@@
In order for the Myers term to be Hermitian, $S_{IJK}$ must be purely imaginary. [^33] Note that if we don’t impose Hermiticity of the Myers term, there is a solution where everything vanishes except the Myers term and $f_{IJK}^{A}$. This constrains
@@
From these we obtain [^34] Note that because of the second line, the fully antisymmetrized part over $I,J,K$ of $f^{I}_{abJK}$ has to vanish. Thus, unlike the previous case there’s no free antisymmetric anti-self-dual part.
@@
[…]
@@ A.2.1 Type I : $SO(6)\to SO(3)\times SO(3)$
The first case is $m_{IJK}\neq 0$. If some component of $m_{IJK}$ is non-vanishing, one can perform $SO(6)$ rotations of $\mathrm{Re}\ m_{IJK}$ such that one sets to 0 all components in $\mathrm{Re}\ m_{1IJ}$ different than $\mathrm{Re\ }m_{123}$ and $\mathrm{Re\ }m_{145}$, and we choose $\mathrm{Re}\ m_{123}\neq 0$ [^35] Note that by the self-duality condition, if $m_{IJK}$ is non-vanishing, some components necessarily have non-vanishing real parts. Without loss of generality, we choose this real part to lie in $m_{123}$. Note that by self-duality of $m_{IJK}$, the imaginary part of $m_{IJK}$ is expressed in terms of the real part of $m_{IJK}$. One can then show by taking suitable components of (128), that all components of $m_{IJK}$ other than $\mathrm{Re}\ m_{123}$ and $\mathrm{Im}\ m_{456}=-\mathrm{Re}m_{123}$ vanish. Namely we obtain (up to $SO(6)$ rotations),
@@
Now taking a linear combination of these two equations [^36] More specifically, $E^{abN}_{JKN}-\frac{i}{3!}E^{abN}_{PQR}\epsilon^{PQRJKN}-F_{JK}^{ab}$, one obtains that the only possibility is $m^{ab}_{I}=0$. For convenience we choose m=1/96m=-1/96. Putting everything together, we found the first solution
[…]
@@
Note that this action is invariant under $SO(2,1)\times SO(3)$ rotations, and under $SU(2)$ transformations acting on the symplectic index, $\psi_{a}\to U_{ab}\psi_{b},U\in SU(2)$. The bosonic symmetry group is thus isomorphic to $SO(4)\times Sl(2,\mathbb{R})$ which forms the bosonic part of the exceptional continuous family of superalgebras $D(2,1,\alpha)$. To ensure that $D(2,1,\alpha)$ is indeed the superalgebra and compute $\alpha$, it is important to compute the anticommutators of two supercharges. To obtain a superalgebra which closes off-shell, we will introduce a set of auxiliary fields $K_{m=1,\ldots,3}$, and consider the action
@@
[…]
where $\nu_{a}^{m}(\epsilon)$ is a spinorial parameter that will be constrained in terms of $\epsilon$ to obtain closure off-shell. This construction is inspired by Berkovits:1993hx; Asano:2012zt. Now we can compute the action of the anticommutator $\{\delta_{\epsilon},\delta_{\eta}\}$ on the relevant matrices exactly as in (Komatsu:2024ydh, App.A.1). Imposing the closure of the algebra with the introduction of the auxiliary fields imposes conditions on $\nu_{m}^{a}$ that restricts it in terms of $\epsilon$. To obtain these constraints, it is sufficient to look at the undeformed model ($\mu=0$). We obtain in particular
@@
[…]
@@
[…]
where the last line also applies to $T^{(1)}$ and $T^{(2)}$. As a consistency check, note that the three coefficients in the anticommutators sum to zero, $\frac{3}{4}+\frac{1}{4}+\frac{1}{2}=0$ as required by Jacobi identities Frappat:1996pb. Up to rescaling the supercharge, this is exactly the $D(1,2,\alpha=\frac{1}{2})$ superalgebra that can be found in the literature Ivanov:2002pc; Frappat:1996pb.